You order fifteen burritos to go from a Mexican restaurant, five with hot peppers and ten without. However, the restaurant forgot to label them. If you pick four burritos at random, find the probability of the given event. (Round your answer to three decimal places.)

Exactly one has hot peppers.
Answer #1

Here x=5 and n=15

So p=x/n=5/15=0.333

Here we see that p=0.333 is same for all, n=4 is constant, events are independent and only two outcomes

So properties of binomial are satisfied

So